A 13-N force moves through a displacement of 42 m, at an angle of 35o to the
displacement. What work is now done by the force? What if the angle is 90 degrees?
1
Expert's answer
2020-11-25T10:53:35-0500

By definition, the work is:


"A = \mathbf{F}\cdot \mathbf{s}"

Where "\mathbf{F}" is vector of the foce and "\mathbf{s}" is vector of the displacement. "\cdot" denotes scalar product. Thus, obtain:


"A = Fs\cos\theta"

where "F = 13N" is the magnitude of the force, "s = 42m" is the magnitude of the displacement, and "\theta=35\degree" is the angle between the work and displacement vectors. Get:

"A = Fs\cos\theta = 13\cdot 42\cdot \cos35\degree \approx 447.3J"

If "\theta =90\degree", then "\cos\theta = 0" and the work is equal to 0 J.

Answer. a) 447.3 J, b) 0 J.